You briefly tap a 0.85 kg toy car initially at rest forward by hand with a force of 10.0 newtons.

The interaction of your hand with the car was .10 seconds (assume zero friction).

If the car was initially at rest, what is the speed of the car immediately after it is tapped?
Answer #1


accelaration a = F/m

from equations of motion


v = u + a*t


u = initial velocity = 0


t = time 10 s


a = acceleration = F/m


v = 0+(F/m)*t

v = (10/0.85)*10


v = 117.65 m/s <<--------answer